40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 4 20:42:53 2020, gautam, Update, Electronics, C1PSL acromag crate is sitting on the floor 
    Reply  Thu Mar 5 11:27:48 2020, gautam, Update, Electronics, C1PSL acromag crate is no longer sitting on the floor 
       Reply  Thu Mar 5 19:51:14 2020, gautam, Update, Electronics, IMC Servo board being tested 
          Reply  Fri Mar 6 01:12:10 2020, gautam, Update, Electronics, IMC Servo IN2 path looks just fine AO_inputTFs_5Mar.pdfsliderCal_5Mar.pdf
Message ID: 15258     Entry time: Fri Mar 6 01:12:10 2020     In reply to: 15257
Author: gautam 
Type: Update 
Category: Electronics 
Subject: IMC Servo IN2 path looks just fine 

It seems like the AO path gain stages on the IMC Servo board work just fine. The weird results I reported earlier were likely a measurement error arising from the fact that I did not disconnect the LEMO IN2 cable while measuring using the BNC IN2 connector, which probably made some parasitic path to ground that was screwing the measurement up. Today, I re-did the measurement with the signal injected at the IN2 BNC, and the TF measured being the ratio of TP3 on the board to a split-off of the SR785 source (T-eed off). Attachments #1, #2 shows the result - the gain deficit from the "expected" value is now consistent with that seen on other sliders.

Note that the signal from the CM board in the LSC rack is sent single-ended over a 2-pin LEMO cable (whose return pin is shorted to ground). But it is received differentially on the IMC Servo board. I took this chance to look for evidence of extra power line noise due to potential ground loops by looking at the IMC error point with various auxiliary cables connected to the board - but got distracted by some excess noise (next elog).

Attachment 1: AO_inputTFs_5Mar.pdf  626 kB  | Hide | Hide all
AO_inputTFs_5Mar.pdf
Attachment 2: sliderCal_5Mar.pdf  110 kB  | Hide | Hide all
sliderCal_5Mar.pdf
ELOG V3.1.3-